OpenStudy (anonymous):

Jackson is playing a video game that has 4 castles to conquer within each level. Jackson has completed 5 levels so far. If x represents the number of remaining levels to complete, which of the following equations can be used to find the total number of castles Jackson will have conquered once he completes the final level?

OpenStudy (anonymous):

y = 4x + 20 y = 4x + 5 y = 20x + 4 y = 5x + 4